# TrimCaching: Parameter-sharing Edge Caching for AI Model Downloading

### v3

@@ Abstract
Index Terms:
Edge AI model caching, edge computing, edge intelligence, 6G, model downloading.
@@ I Introduction
In this paper, we address a novel AI model placement problem by exploiting parameter sharing to improve model storage efficiency at the network edge. A broad range of AI models, such as convolutional neural networks (CNNs) or LLMs, can share a significant proportion of parameters, as they can be derived from the same pre-trained models [18, 19]. For example, bottom-layer freezing is a classic technique in transfer learning and multi-task learning, since the bottom layers in deep neural networks, such as convolution layers, usually share common knowledge reusable for different downstream tasks [20, 21, 22, 23]. In the era of LLMs, parameter-efficient fine-tuning (PEFT), such as LoRA [24], has emerged as an effective approach to adapt foundation models for downstream applications by freezing a large proportion of parameters, thereby saving computing and memory resources. For example, LoRA freezes all parameters in the pre-trained LLMs and introduces only a small number of trainable parameters for fine-tuning, typically less than 1% of the total, implying over 99% of the parameters remain unchanged during fine-tuning. As shown in Fig. 1, the inference accuracy of a fine-tuned ResNet50 degrades only slightly as the number of bottom layers frozen from the pre-trained model increases. Even when the first 90% of trainable layers, up to layer 97, are frozen, the average accuracy degradation is only about 4.7%, compared with the full-layer fine-tuning. As for LLMs, Fig. 2 illustrates that the BLEU scores of various fine-tuned GPT-2 medium models exhibit slight variation across different frozen parameter ratios. The fine-tuned models maintain satisfactory performance, even when up to 99.97% of the parameters are frozen from the pre-trained model. In a nutshell, downstream models can share a significant proportion of parameters from the same pre-trained model, inherently making storage more efficient.
@@ III-A Network Scenario
We consider a multi-edge multi-user scenario in wireless edge networks, as illustrated in Fig. 3. Let $\mathcal{M}=\left\{1,\dots,m\dots,M\right\}$ denote the set of interconnected wireless edge servers (e.g., base stations), with a set of users ${\mathcal{K}}=\left\{1,\dots,k,\dots,K\right\}$ distributed across their coverage area. To support fast AI model downloading services for inference, these edge servers collectively cache a set of AI models ${\mathcal{I}}=\left\{1,\dots,i,\dots,I\right\}$. User $k$ requests downloading model $i$ for performing local inference with a probability $p_{k,i}$ [^1] In practice, $p_{k,i}$ can be obtained by monitoring and analyzing user historical traffic [47].. We use $\bar{T}_{k,i}$ to denote the E2E latency requirement of user $k$ for consuming model $i$, which consists of model downloading and local inference latency. The frequently used notations are summarized in Table I.
@@ III-D E2E latency
  • Downloading from non-associated edge servers: If none of the edge servers in $\mathcal{M}_{k}$ caches model $i$, the model will be fetched from another edge server $m$ in $\mathcal{M}\setminus\mathcal{M}_{k}$ where the model is cached. Specifically, the model is delivered from edge server $m$ to an edge server $m^{\prime}$ in $\mathcal{M}_{k}$ and then to user $k$. A cache hit occurs if the E2E latency (including edge-to-edge, edge-to-user, and on-device inference latency) meets latency requirement $\bar{T}_{k,i}$.
@@ IV-A Problem Formulation
Note that our formulation ignores user mobility because the problem is solved based on a “snapshot” of user locations. This is commonly adopted in model placement schemes [30]. In practice, our algorithm can conduct model placement decisions by solving the above problem, then re-initiate model placement when the performance degrades to a certain threshold. Our simulation results will show that our algorithm is resilient to user mobility over time, thus eliminating the need for frequent model replacement that would consume backbone bandwidth.
@@ V-B TrimCaching Spec Algorithm
where
@@ V-C3 Optimal model caching of models in $\mathcal{I}_{\mathcal{N}}$ on an edge server
  • $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}\geq u\left(m,e_{\mathcal{N}}\right)$, implying that the number of cache hits of the $e_{\mathcal{N}}$-th model is no greater than the remaining number of cache hits to be satisfied.
  • $\mathcal{T}\left(e_{\mathcal{N}},\dot{w}_{\mathcal{N}}-\frac{\dot{u}\left(m,e_{\mathcal{N}}\right)}{\dot{\delta}_{m,\mathcal{N}}}\right)+D\left(e_{\mathcal{N}}\right)<\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}\right)$, representing that the number of cache hits $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}$ is achieved by model $e_{\mathcal{N}}$, contributing $\dot{u}\left(m,e_{\mathcal{N}}\right)$, and a subset of the first $e_{\mathcal{N}}-1$ models, contributing $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}-\dot{u}\left(m,e_{\mathcal{N}}\right)$. Besides, the corresponding total data size, $\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}-\frac{\dot{u}\left(m,e_{\mathcal{N}}\right)}{\dot{\delta}_{m,\mathcal{N}}}\right)+D\left(e_{\mathcal{N}}\right)$, is less than $\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}\right)$.
@@
Input: $m$, $\mathcal{I}_{\mathcal{N}}$, $\dot{\delta}_{m,\mathcal{N}}$, $u\left(m,e_{\mathcal{N}}\right)$, and $\mathcal{T}\left(e_{\mathcal{N}},\dot{w}_{\mathcal{N}}\right)$.
@@
13    for $\dot{w}_{\mathcal{N}}\in\left\{1,\dots,\dot{W}_{\mathcal{N}}\right\}$ do
@@ VII-A Simulation Setup
  • General case: In this case, we design parameter sharing in the ResNet-based model library. To support arbitrary parameter sharing among models in the library (shared parameter blocks do not simply come from a small set of pretrained models, as in the special case, and the number of shared blocks increases as the library grows), we adopt a two-stage fine-tuning strategy. In the first stage, we select multiple superclasses from CIFAR100. For each selected superclass, all parameters of the pre-trained ResNet-18, ResNet-34, and ResNet-50 are fine-tuned using data from all its classes, resulting in one distinct model per structure for each superclass. In the second stage, each first-stage fine-tuned model is further fine-tuned on datasets from classes within a new superclass that is semantically similar to the one used in the first stage. The pairing of superclasses between stage 1 and stage 2 is provided in Table II. Moreover, the second-stage fine-tuning is conducted using bottom-layer freezing, without restricting the number of frozen layers, where each first-stage model is fine-tuned separately on the data of each class, resulting in a total of 45 models.
Moreover, the on-device inference latency for the ResNet-based and the GPT-2-based model libraries ranges from 1 ms to 5 ms and 40 ms to 200 ms, respectively [10]. The request probabilities of each user for the models in each model library follow a Zipf distribution [65].
@@ VII-F Impact of Request Probability Uncertainty on Cache Hit Ratio
This subsection investigates the effect of the uncertainty of model request probabilities on the cache hit ratio, as the estimated request probabilities used for model caching decision determination may deviate from true value $p_{k,i}$. Following prior work, we model the estimated request probability of user $k$ for model $i$, denoted by $p^{\prime}_{k,i}$, as a beta distribution with mean $p_{k,i}$ [66, 67]. In our experiments, the model caching decisions are first determined based on $p^{\prime}_{k,i}$, while the cache hit ratio is evaluated through Monte Carlo experiments, where each user generates a specific model request according to $p_{k,i}$ in each Monte Carlo experiment. We analyze the cache hit ratio in terms of the coefficient of variation of $p^{\prime}_{k,i}$.
Fig. 12 demonstrates that the cache hit ratio of the proposed algorithms decreases slightly as the coefficient of variation of $p^{\prime}_{k,i}$ increases. Fig. 12(a) shows that the cache hit ratio of the TrimCaching Spec and TrimCaching Gen algorithms degrades by 2.20% and 2.78%, respectively, as the coefficient of variation of $p^{\prime}_{k,i}$ increases from 0 to 0.5 in the special case. Besides, in the general case, the cache hit ratio of the TrimCaching Gen algorithms drops by 1.82% over the same range in Fig. 12(b). These results demonstrate that the proposed algorithms remain robust with only minor performance degradation even when the estimated request probabilities deviate from their true values substantially.
@@
Figure 12: Cache hit ratio versus the coefficient of variation of estimated request probabilities using the ResNet-based model library, where $Q=$ 1 GB, $K=$ 30, $M=$ 10.
@@ VII-G Comparison with Online Placement Strategies
In this section, we compare the proposed algorithms with two classical online content placement strategies, including the Least Recently Used (LRU) and the Least Frequently Used (LFU) strategies, under both the special and general cases. The experiments are conducted in a dynamic setting, where each user generates one specific model request according to $p_{k,i}$ at the beginning of each 10-second time slot. The placement decisions of the TrimCaching Spec and TrimCaching Gen algorithms are determined once at time 0 based on $p_{k,i}$ and remain unchanged throughout the experiment. In contrast, both LRU and LFU update their placements dynamically according to the user requests at the end of each time slot. Under LRU, models that are least recently requested are removed from the cache, while under LFU, those with the least request frequency are removed [68, 69]. We consider parameter sharing in both LRU and LFU implementations. The cache hit ratio is recorded at the start of each time slot to represent the performance for that slot.
@@
Figure 13: Cache hit ratio comparisons with online placement strategies using the ResNet-based model library, where $Q$, $K$, and $M$ follow Fig. 12.
@@ Appendix D Proof of Proposition 4
Let ${\tilde{\bf{X}}}=\bigcup\limits_{m\in\mathcal{M}}{\tilde{\bf{X}}_{m}}$ denote the set of placement decisions made by $\bf{X}$$^{*}$ but not by $\hat{{\bf{X}}}$, i.e., $\tilde{x}_{m,i}\in\ {\tilde{\bf{X}}}$ satisfies
@@
We begin by assuming that $U\left({\tilde{\bf{X}}}\right)=U\left(\hat{{\bf{X}}}\right)+\Delta$ holds for some $\Delta>0$. Next, according to Proposition 3, we have $U\left({\tilde{\bf{X}}}\right)=U\left(\bigcup\limits_{m\in\mathcal{M}}{\tilde{\bf{X}}}_{m}\right)=\sum\limits_{m\in\mathcal{M}}\hat{U}_{m}\left({\tilde{\bf{X}}}_{m}\right)$, where $\hat{U}_{m}\left(\cdot\right)$ is defined in (8). Although Proposition 3 is for $\hat{{\bf{X}}}$, the decomposition applies to $\tilde{\bf{X}}$ by the same reasoning. This implies that
@@
$>U\left(\hat{{\bf{X}}}\right)=\sum\limits_{m\in\mathcal{M}}\hat{U}_{m}\left(\hat{{\bf{X}}}_{m}\right).$

### v4

@@ I Introduction
In this paper, we address a novel AI model placement problem by exploiting parameter sharing to improve model storage efficiency at the network edge. A broad range of AI models, such as convolutional neural networks (CNNs) or LLMs, can share a significant proportion of parameters, as they can be derived from the same pre-trained models [18, 19]. For example, bottom-layer freezing is a classic technique in transfer learning and multi-task learning, since the bottom layers in deep neural networks, such as convolution layers, usually share common knowledge reusable for different downstream tasks [20, 21, 22, 23]. In the era of LLMs, parameter-efficient fine-tuning (PEFT), such as LoRA [24], has emerged as an effective approach to adapt foundation models for downstream applications by freezing a large proportion of parameters, thereby saving computing and memory resources. For example, LoRA freezes all parameters in the pre-trained LLMs and introduces only a small number of trainable parameters for fine-tuning, typically less than 1% of the total, implying over 99% of the parameters remain unchanged during fine-tuning. As shown in Fig. 1, the inference accuracy of a fine-tuned ResNet-50 degrades only slightly as the number of bottom layers frozen from the pre-trained model increases. Even when the first 90% of trainable layers, up to layer 97, are frozen, the average accuracy degradation is only about 4.7%, compared with the full-layer fine-tuning. As for LLMs, Fig. 2 illustrates that the BLEU scores of various fine-tuned GPT-2 medium models exhibit slight variation across different frozen parameter ratios. The fine-tuned models maintain satisfactory performance, even when up to 99.97% of the parameters are frozen from the pre-trained model. In a nutshell, downstream models can share a significant proportion of parameters from the same pre-trained model, inherently making storage more efficient.
@@ III-A Network Scenario
We consider a multi-edge multi-user scenario in wireless edge networks, as illustrated in Fig. 3. Let $\mathcal{M}=\left\{1,\dots,m\dots,M\right\}$ denote the set of interconnected wireless edge servers (e.g., base stations), with a set of users ${\mathcal{K}}=\left\{1,\dots,k,\dots,K\right\}$ distributed across their coverage areas. To support fast AI model downloading services for inference, these edge servers collectively cache a set of AI models ${\mathcal{I}}=\left\{1,\dots,i,\dots,I\right\}$. User $k$ requests to download model $i$ for performing local inference with probability $p_{k,i}$ [^1] In practice, $p_{k,i}$ can be obtained by monitoring and analyzing user historical traffic [47].. We use $\bar{T}_{k,i}$ to denote the E2E latency requirement of user $k$ for consuming model $i$, which consists of model downloading and local inference latency. The frequently used notations are summarized in Table I.
@@ III-D E2E latency
  • Downloading from non-associated edge servers: If none of the edge servers in $\mathcal{M}_{k}$ caches model $i$, the model will be fetched from another edge server $m$ in $\mathcal{M}\setminus\mathcal{M}_{k}$ where the model is cached. Specifically, the model is delivered from edge server $m$ to an edge server $m^{\prime}$ in $\mathcal{M}_{k}$ and then to user $k$. A cache hit occurs if the E2E latency (including edge-to-edge, edge-to-user, and on-device inference latency) meets the latency requirement $\bar{T}_{k,i}$.
@@ IV-A Problem Formulation
Note that our formulation ignores user mobility because the problem is solved based on a “snapshot” of user locations. This is commonly adopted in model placement schemes [30]. In practice, our algorithm can make model placement decisions by solving the above problem, then re-initiate model placement when the performance degrades to a certain threshold. Our simulation results will show that our algorithm is resilient to user mobility over time, thus eliminating the need for frequent model replacement that would consume backbone bandwidth.
@@ V-B TrimCaching Spec Algorithm
where [^2] For notation simplicity, $\hat{U}_{m}\left(\hat{{\bf{X}}}_{m}\right)$ is used as shorthand for $\hat{U}_{m}\left(\hat{{\bf{X}}}_{m}\middle|\bigcup\limits_{m^{\prime}=1}^{m-1}\hat{{\bf{X}}}_{m^{\prime}}\right)$.
@@ V-C3 Optimal model caching of models in $\mathcal{I}_{\mathcal{N}}$ on an edge server
  • $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}\geq\dot{u}\left(m,e_{\mathcal{N}}\right)$, implying that the number of cache hits of the $e_{\mathcal{N}}$-th model is no greater than the remaining number of cache hits to be satisfied.
  • $\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}-\frac{\dot{u}\left(m,e_{\mathcal{N}}\right)}{\dot{\delta}_{m,\mathcal{N}}}\right)+D\left(e_{\mathcal{N}}\right)<\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}\right)$, representing that the number of cache hits $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}$ is achieved by model $e_{\mathcal{N}}$, contributing $\dot{u}\left(m,e_{\mathcal{N}}\right)$, and a subset of the first $e_{\mathcal{N}}-1$ models, contributing $\dot{w}_{\mathcal{N}}\dot{\delta}_{m,\mathcal{N}}-\dot{u}\left(m,e_{\mathcal{N}}\right)$. Besides, the corresponding total data size, $\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}-\frac{\dot{u}\left(m,e_{\mathcal{N}}\right)}{\dot{\delta}_{m,\mathcal{N}}}\right)+D\left(e_{\mathcal{N}}\right)$, is less than $\mathcal{T}\left(e_{\mathcal{N}}-1,\dot{w}_{\mathcal{N}}\right)$.
@@
Input: $m$, $\mathcal{I}_{\mathcal{N}}$, $\dot{\delta}_{m,\mathcal{N}}$, $\dot{u}\left(m,e_{\mathcal{N}}\right)$, and $\mathcal{T}\left(e_{\mathcal{N}},\dot{w}_{\mathcal{N}}\right)$.
@@
13    for $\dot{w}_{\mathcal{N}}\in\left\{1,\dots,\dot{W}_{m,\mathcal{N}}\right\}$ do
@@ VII-A Simulation Setup
  • General case: In this case, we design parameter sharing in the ResNet-based model library. To support arbitrary parameter sharing among models in the library (shared parameter blocks do not simply come from a small set of pre-trained models, as in the special case, and the number of shared blocks increases as the library grows), we adopt a two-stage fine-tuning strategy. In the first stage, we select multiple superclasses from CIFAR100. For each selected superclass, all parameters of the pre-trained ResNet-18, ResNet-34, and ResNet-50 are fine-tuned using data from all its classes, resulting in one distinct model per structure for each superclass. In the second stage, each first-stage fine-tuned model is further fine-tuned on datasets from classes within a new superclass that is semantically similar to the one used in the first stage. The pairing of superclasses between stage 1 and stage 2 is provided in Table II. Moreover, the second-stage fine-tuning is conducted using bottom-layer freezing, without restricting the number of frozen layers, where each first-stage model is fine-tuned separately on the data of each class, resulting in a total of 45 models.
Moreover, the on-device inference latency for the ResNet-based and the GPT-2-based model libraries ranges from 1 ms to 5 ms and from 40 ms to 200 ms, respectively [10]. The request probabilities of each user for the models in each model library follow a Zipf distribution [65].
@@ VII-F Impact of Request Probability Uncertainty on Cache Hit Ratio
This subsection investigates the effect of the uncertainty of model request probabilities on the cache hit ratio, as the estimated request probabilities used for model caching decision determination may deviate from the true value $p_{k,i}$. Following prior work, we model the estimated request probability of user $k$ for model $i$, denoted by $p^{\prime}_{k,i}$, as a Beta distribution with mean $p_{k,i}$ [66, 67]. In our experiments, the model caching decisions are first determined based on $p^{\prime}_{k,i}$, while the cache hit ratio is evaluated through Monte Carlo experiments, where each user generates a specific model request according to $p_{k,i}$ in each Monte Carlo experiment. We analyze the cache hit ratio in terms of the coefficient of variation of $p^{\prime}_{k,i}$.
Fig. 12 demonstrates that the cache hit ratio of the proposed algorithms decreases slightly as the coefficient of variation of $p^{\prime}_{k,i}$ increases. Fig. 12(a) shows that the cache hit ratio of the TrimCaching Spec and TrimCaching Gen algorithms degrades by 2.20% and 2.78%, respectively, as the coefficient of variation of $p^{\prime}_{k,i}$ increases from 0 to 0.5 in the special case. Besides, in the general case, the cache hit ratio of the TrimCaching Gen algorithm drops by 1.82% over the same range in Fig. 12(b). These results demonstrate that the proposed algorithms remain robust with only minor performance degradation even when the estimated request probabilities deviate from their true values substantially.
@@
Figure 12: Cache hit ratio versus the coefficient of variation of estimated request probabilities using the ResNet-based model library, where $Q=$ 1 GB, $K=$ 30, and $M=$ 10.
@@ VII-G Comparison with Online Placement Strategies
In this subsection, we compare the proposed algorithms with two classical online content placement strategies, including the Least Recently Used (LRU) and the Least Frequently Used (LFU) strategies, under both the special and general cases. The experiments are conducted in a dynamic setting, where each user generates one specific model request according to $p_{k,i}$ at the beginning of each 10-second time slot. The placement decisions of the TrimCaching Spec and TrimCaching Gen algorithms are determined once at time 0 based on $p_{k,i}$ and remain unchanged throughout the experiment. In contrast, both LRU and LFU update their placements dynamically according to the user requests at the end of each time slot. Under LRU, models that are least recently requested are removed from the cache, while under LFU, those with the least request frequency are removed [68, 69]. We consider parameter sharing in both LRU and LFU implementations. The cache hit ratio is recorded at the start of each time slot to represent the performance for that slot.
@@
Figure 13: Cache hit ratio comparisons with online placement strategies using the ResNet-based model library, where $Q$, $K$, and $M$ follow those in Fig. 12.
@@ Appendix D Proof of Proposition 4
Let $\tilde{\bf{X}}_{m}$ denote the set of placement decisions for edge server $m$ made by ${\bf{X}}_{m}^{*}$ but not by $\hat{{\bf{X}}}_{m}$, i.e., $\tilde{x}_{m,i}\in\ {\tilde{\bf{X}}}_{m}$ satisfies
@@
where ${\bf{X}}_{m}^{*}$ is the component of the optimal solution $\bf{X}$$^{*}$ associated with edge server $m$, with ${\bf{X}}^{*}=\bigcup\limits_{m\in\mathcal{M}}{\bf{X}}_{m}^{*}$, $x^{*}_{m,i}\in{\bf{X}}_{m}^{*}$, and $\hat{x}_{m,i}\in\hat{{\bf{X}}}_{m}$. Moreover, based on the definition in (8), $\hat{U}_{m}\left(\tilde{{\bf{X}}}_{m}\right)$ is given in (26) at the bottom of this page, which is used as shorthand for $\hat{U}_{m}\left(\tilde{{\bf{X}}}_{m}\middle|\bigcup\limits_{m^{\prime}=1}^{m-1}\hat{{\bf{X}}}_{m^{\prime}}\right)$. Note that ${\tilde{\bf{X}}}_{m}$ and $\hat{U}_{m}\left(\tilde{{\bf{X}}}_{m}\right)$ are introduced solely as auxiliary variables to analyze the theoretical performance gap between $\hat{\bf{X}}$ and $\bf{X}$$^{*}$ after all $\hat{\bf{X}}_{m}$ have been obtained.
$\hat{U}_{m}\left({\tilde{\bf{X}}}_{m}\right)$ $=\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}\tilde{x}_{m,i}\mathbb{I}_{\left\{T_{m,k,i}\leq\bar{T}_{k,i}\right\}}\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}\mathbb{I}_{\left\{T_{m^{\prime},k,i}\leq\bar{T}_{k,i}\right\}}\right)}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}}$ (26)
$=\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}\tilde{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}{p}_{k,i}}.$
@@
$\hat{U}_{m}\left({\tilde{\bf{X}}}_{m}\right)\leq\hat{U}_{m}\left(\hat{{\bf{X}}}_{m}\right),\ \forall m\in\mathcal{M}.$ (27)
Furthermore, summing over all $m$ and invoking Proposition 3, we obtain
@@
Second, we show that
$U\left({\bf{X}}^{*}\right)\leq U\left(\hat{{\bf{X}}}\right)+\sum\limits_{m\in\mathcal{M}}\hat{U}_{m}\left({\tilde{\bf{X}}}_{m}\right).$ (29)
Based on (4), for each model request $r_{k,i}$ of user $k$ for model $i$, the cache hit indicator under $\bf{X}$ is $1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)$, which equals 1 if $r_{k,i}$ is served by $\bf{X}$.
To prove (29), we begin by claiming that for any $r_{k,i}$, the following inequality holds.
$1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)$ (30)
$\leq 1-\prod\limits_{m\in\mathcal{M}}\left(1-\hat{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\right)$
$+\sum\limits_{m\in\mathcal{M}}\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right).$
We verify this inequality by considering two cases.
(i) If $r_{k,i}$ is not served by $\bf{X}^{*}$, then $1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=0$. Since the right-hand side of (30) is nonnegative, (30) holds in this case.
(ii) If $r_{k,i}$ is served by $\bf{X}^{*}$, then $1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)$ on the left-hand side of (30) equals 1. We further consider two sub-cases. (ii-a) If $r_{k,i}$ is also served by $\hat{\bf{X}}$, then $1-\prod\limits_{m\in\mathcal{M}}\left(1-\hat{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=1$. Therefore, $1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=1-\prod\limits_{m\in\mathcal{M}}\left(1-\hat{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)$, and (30) holds in this sub-case. (ii-b) If $r_{k,i}$ is not served by $\hat{\bf{X}}$, then $1-\prod\limits_{m\in\mathcal{M}}\left(1-\hat{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=0$, implying $\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)=1$ for all $m\in\mathcal{M}$. Since $r_{k,i}$ is served by $\bf{X}$$^{*}$ but not by $\hat{\bf{X}}$, $r_{k,i}$ is served by $\bigcup\limits_{m\in\mathcal{M}}{\tilde{\bf{X}}}_{m}$, which implies $1-\prod\limits_{m\in\mathcal{M}}\left(1-\tilde{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=1$. Therefore, there exists at least one $m\in\mathcal{M}$ such that $\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)=1$, and consequently $\sum\limits_{m\in\mathcal{M}}\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\geq 1$. Combining this with the fact that, in this sub-case, $\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)=1$ for all $m\in\mathcal{M}$, we obtain $\sum\limits_{m\in\mathcal{M}}\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)\geq 1$. Since $1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)=1$ in this sub-case, (30) holds in this sub-case. Therefore, (30) holds for all $r_{k,i}$.
Multiplying both sides of (30) by $p_{k,i}$, summing over all $r_{k,i}$, and normalizing by $\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}$, we obtain (31) and (32), which are given at the bottom of the next page. Since the second term on the right-hand side of (32) equals $\sum\limits_{m\in\mathcal{M}}\hat{U}_{m}\left({\tilde{\bf{X}}}_{m}\right)$, (29) holds.
$\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\left[1-\prod\limits_{m\in\mathcal{M}}\left(1-x_{m,i}^{*}{\mathbb{I}}_{1}\left(m,k,i\right)\right)\right]}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}}$ $\leq\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\left[1-\prod\limits_{m\in\mathcal{M}}\left(1-\hat{x}_{m,i}{\mathbb{I}}_{1}\left(m,k,i\right)\right)\right]}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}}$ (31)
$+\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\left[\sum\limits_{m\in\mathcal{M}}\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)\right]}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}}.$
$U\left({\bf{X}}^{*}\right)\leq U\left(\hat{{\bf{X}}}\right)+\sum\limits_{m\in\mathcal{M}}\frac{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}\tilde{x}_{m,i}\mathbb{I}_{1}\left(m,k,i\right)\prod\limits_{m^{\prime}=1}^{m-1}\left(1-\hat{x}_{m^{\prime},i}{\mathbb{I}}_{1}\left(m^{\prime},k,i\right)\right)}{\sum\limits_{k\in\mathcal{K}}\sum\limits_{i\in\mathcal{I}}p_{k,i}}.$ (32)
Finally, substituting (28) into (29) yields